A chemist plans to mix 40 ml of a 5% alcohol solution with some 11% alcohol solution to make a 7% solution. How much of the 11% solution should be mixed with the 5% alcohol solution to obtain the 7% alcohol solution?
Answer by checkley77(12844)

.11X+.05*40=.07(40+X)
.11X+2=2.8+.07X
.11X-.07X=2.8-2
.04X=.8
X=.8/.04
X=20 ML. OF 11% ALCOHOL IS USED.
PROOF:
.11*20+2=.07(40+20)
2.2+2=.07*60
4.2=4.2